OpenStudy (anonymous):

2^(7x)=23 help please

OpenStudy (dumbcow):

rewrite it in log form \[7x = \log_{2} 23 \] \[x = \frac{\log 23}{7\log 2}\]

OpenStudy (tommo_lcfc):

Use logarithms. x = 0.646 (3 dp) Follow @dumbcow's method